OpenStudy (anonymous):

If the common difference in an arithmetic sequence is 4, and the 20th term is 36, what is the first term?

OpenStudy (solomonzelman):

Given:\(\large\color{black}{ \\[0.6 em] }\) \(\large\color{black}{ \displaystyle {\rm \color{blue}{d}}=4 }\)\(\large\color{black}{ \\[0.6 em] }\) \(\large\color{black}{ \displaystyle \color{blue}{a_{20}}=36 }\)

OpenStudy (solomonzelman):

Formula\(\large\color{black}{ \\[0.6 em] }\) \(\Large\color{black}{ \displaystyle a_{{\color{red}{\rm n}}}=\left(a_{\color{green}{\rm 1}}\right)+{\rm \color{blue}{d}}\left({\color{red}{\rm n}}-{\color{green}{\rm 1}}\right) }\)\(\large\color{black}{ \\[0.6 em] }\) I labelled the corresponding parts in colors.